i must find the area of triangle ABC when a=40ft b=25ft and angle c=30 degrees please help i dont understand this at all
Answer by stanbon(75887)

i must find the area of triangle ABC when a=40ft b=25ft and angle c=30 degrees please help i dont understand this at all
---------------
Chances are you are talking about a right triangle.
----
If that is true:
Area = a*b*sin(C) = 40*25*(1/2) = 20*25 = 500 sq. ft
========================================================
Cheers,
Stan H.